OpenStudy (anonymous):

help please! Solve x^2-4=5 by graphing the related function. 1)there are two solutions 1 and -1 2)there is one solution 1 3)there are no real number solutions. 4)there are two solutions 3 and -3

OpenStudy (anonymous):

x^2-4=5 x^2-4-5=0 x^2-9=0 (x)^2-(3)^2 = 0 (x+3)(x-3)=0 so you have two roots/solutions...... -3 and 3

OpenStudy (campbell_st):

well if you graph y = x^2 - 4 and then graph y = 5 the 2 points of intersection will give the solutions...